# Anisotropy of quadratic forms over global fields of characteristic $\neq$ 2 is Diophantine

### v2

@@ 2.2. Global class field theory
In this section, we recall some standard results from global class field theory. We begin by introducing the notion of a modulus, which generalizes the integer $m$ in classical modular arithmetic.
@@
$I_{\mathfrak{m}}:=\langle\mathfrak{p}\subseteq\mathcal{O}_{K}:\mathfrak{m}(\mathfrak{p})=0\rangle$
@@
$(\alpha):=\alpha\mathcal{O}_{K}\;\in\;I_{K}.$
@@
$P_{\mathfrak{m}}:=i(K_{\mathfrak{m},1})\;\leq\;I_{\mathfrak{m}}.$
@@
$C_{\mathfrak{m}}:=I_{\mathfrak{m}}/P_{\mathfrak{m}}.$
@@
$\psi_{L/K}(\mathfrak{a})\;:=\;\prod(\mathfrak{p}_{i},L/K)^{e_{i}}.$
@@
where the first product runs over all finite primes $\mathfrak{P}$ of $L$ above $\mathfrak{p}$, and the second runs over all real embeddings $\Omega$ of $L$ extending the real embedding $\omega$ of $K$.
@@ 3.1. Notations and Some Examples
$S(A/K):=\{\textup{Trd}(x):x\in A,\ \textup{Nrd}(x)=1\},\qquad T(A/K):=S(A/K)-S(A/K),$
@@
$H_{a,b}:=K\oplus\alpha K\oplus\beta K\oplus\alpha\beta K,$
where $\alpha^{2}=a$, $\beta^{2}=b$, $\alpha\beta=-\beta\alpha$.
@@
$\mathcal{O}_{\mathfrak{p}}:=(\mathcal{O}_{K})_{\mathfrak{p}}=R_{\mathfrak{p}}\cap K$
@@
If $A/K$ splits at all real places of $K$ (which is always the case if $K$ is a global function field or $a,b\in K^{\times}$ are totally positive if $K$ is a number field), then
@@
$T_{a,b}$ is a diophantine set over $K$ by definition, and this is the basis to prove that many other sets are diophantine in [EM18],[Dit18] and [Par13]. Actually, $T_{a,b}$ is uniformly existential defined in $a$ and $b$ since we have the following formula for $S_{a,b}$
$S_{a,b}=\{2x_{1}\in K:\exists x_{2},x_{3},x_{4}\textup{ with }x_{1}^{2}-ax_{1}^{2}-bx_{3}^{2}+abx_{4}^{2}=1\}$
as well as the definition $T_{a,b}=S_{a,b}+S_{a,b}$.
The following definition is given in Definition 9 of [Koe16] for $K=\mathbb{Q}$, and also given after the proof of Lemma 3.8 in [Par13] as well as after the proof of Lemma 3.3 in [EM18].
@@
Let $L:=K(\sqrt{a},\sqrt{b})/K$ be an abelian extension with Galois group isomorphic to $(\mathbb{Z}/2\mathbb{Z})^{2}$ and $\psi_{L/K}:I_{\mathfrak{m}}\rightarrow\textup{Gal}(L/K)$ be the global Artin map with the admissible modulus $\mathfrak{m}$. We define the following notations:
  • $\mathbb{P}(p):=\{\mathfrak{p}\in S_{K}^{f}:v_{\mathfrak{p}}(p)\text{ is odd}\}$.
  • For $(i,j)\in\textup{Gal}(L/K)$, $i,j\in\{\pm 1\}$, we define $\mathbb{P}^{(i,j)}=\{\mathfrak{p}\in S_{K}^{f}:\mathfrak{p}\in I_{m}\text{ and }\psi_{L/K}(\mathfrak{p})=(i,j)\}$.
  • $\mathbb{P}^{(i,j)}(p):=\mathbb{P}(p)\cap\mathbb{P}^{(i,j)}$.
Here, $(1,1)$ is the identity in the Galois group. (1,1)(-1,1) is the element which sends $\sqrt{a}$ to $\sqrt{a}$ and fixes $\sqrt{b}$. (1,1)(1,-1) is the element which sends $\sqrt{b}$ to $\sqrt{b}$ and fixes $\sqrt{a}$. Finally, (1,1)(-1,-1) is the element sends $\sqrt{a}$ to $\sqrt{a}$ and $\sqrt{b}$ to $\sqrt{b}$.
@@
(1) If $K$ is a number field, then we can find $a,b\in K^{\times}$ with disjoint support and $\textup{Gal}(K(\sqrt{a},\sqrt{b})/K)\cong(\mathbb{Z}/2\mathbb{Z})^{2}$ such that for any $p\in K^{\times}$ co-prime to the admissible modulus $\mathfrak{m}$, we have
@@
(2) If $K$ is a global function field with $\textup{char}(K)\neq 2$, then we can find $a,b,c,d\in K^{\times}$ with $\textup{Gal}(K(\sqrt{a},\sqrt{b})/K)=(\mathbb{Z}/2\mathbb{Z})^{2}$ such that for any $p\in K^{\times}$ coprime to the the admissible modulus $\mathfrak{m}$, we have
@@
First we explain (1), this is the case proved in [Par13]. From Lemma 3.9 of [Par13] we know that for any $p$ coprime to $\mathfrak{m}$, the sets at the left-hand and right-hand sides of each equation differ at most by the primes dividing the modulus. And then Lemma 3.19 and Lemma 3.20 in [Par13] give a carefully chosen $a$ and $b$ such that the two sets are equal.
The case (2) is Corollary 3.11 in [EM18].
The support of $a$ and $b$ are disjoint by Lemma 3.19 (4) of [Par13] and Lemma 3.8 (4) of [EM18]. ∎
From now on, unless otherwise stated, our parameters $a,b$ or $a,b,c,d$ are the same as in Lemma 3.4, depending on the case of number fields or function fields.
@@
Let $K$ be a global field with $\textup{char}(K)\neq 2$, and keep the constants $a,b,c,d$ as above. We define the ring $R_{p}^{\sigma}$ for $\sigma\in\textup{Gal}(K(\sqrt{a},\sqrt{b})/K)$ as follows:
(a): If $K$ is a number field, then we define
  • $R_{p}^{(-1,-1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{a,p}\cap\Delta_{b,p}}\mathcal{O}_{\mathfrak{p}}$,
  • $R_{p}^{(1,-1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{ab,p}\cap\Delta_{b,p}}\mathcal{O}_{\mathfrak{p}}$,
  • $R_{p}^{(-1,1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{a,p}\cap\Delta_{ab,p}}\mathcal{O}_{\mathfrak{p}}$,
  • $R_{p,q}^{(1,1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{ap,q}\cap\Delta_{bp,q}}\mathcal{O}_{\mathfrak{p}}$.
(b): If $K$ is a global function field with $\textup{char}(K)\neq 2$, then we define
  • $R_{p}^{(-1,-1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{a,p}\cap\Delta_{b,p}}\mathcal{O}_{\mathfrak{p}}$,
  • $R_{p}^{(1,-1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{ab,p}\cap\Delta_{b,p}\cap\Delta_{a,cp}}\mathcal{O}_{\mathfrak{p}}$,
  • $R_{p}^{(-1,1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{a,p}\cap\Delta_{ab,p}\cap\Delta_{b,dp}}\mathcal{O}_{\mathfrak{p}}$,
  • $R_{p,q}^{(1,1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{ap,q}\cap\Delta_{bp,q}}\mathcal{O}_{\mathfrak{p}}$.
For $K=\mathbb{Q}$, the above definition of $R_{p}^{\sigma}$ and $R_{p,q}^{(1,1)}$ is given as Definition 7 in [Koe16]. For the generalization to number fields, it is Definition 3.10 of [Par13]. For global fields, it is Definition 3.12 of [EM18].
@@
Let $\Delta$ be a finite subset of $S_{K}^{f}$. Recall that $J(R)$ is the Jacobson radical of $R$. If $R=\bigcap\limits_{\mathfrak{p}\in\Delta}\mathcal{O}_{\mathfrak{p}}$, then $J(R)=\bigcap\limits_{\mathfrak{p}\in\Delta}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}.$
@@
Let $x\in R$, the inverse $x^{-1}\in K$ lies in $R$ if and only if $v_{\mathfrak{p}}(x^{-1})\geq 0,\forall\mathfrak{p}\in\Delta$. Combining this with $x\in R$ we get $v_{\mathfrak{p}}(x)=0,\forall\mathfrak{p}\in\Delta$. Hence the set of invertible elements in $R$ is
$\{x\in R:v_{\mathfrak{p}}(x)=0,\forall\mathfrak{p}\in\Delta\}.$
Suppose that $x\in\bigcap\limits_{\mathfrak{p}\in\Delta}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$, then $v_{\mathfrak{p}}(1+yx)=0,\forall y\in R,\forall\mathfrak{p}\in\Delta$, hence $x\in J(R)$.
Suppose that $x\in J(R)$, we claim that $v_{\mathfrak{p}}(x)>0$ for all $\mathfrak{p}\in\Delta$, hence $x\in\bigcap\limits_{\mathfrak{p}\in\Delta}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}.$
If there exists $\mathfrak{p}^{\prime}\in\Delta$ such that $v_{\mathfrak{p}^{\prime}}(x)=0$, then $x^{-1}\in\mathcal{O}_{\mathfrak{p}^{\prime}}$, then by Chinese remainder theorem there exists $y\in R$ with $v_{\mathfrak{p}^{\prime}}(y-x^{-1})\geq 1$. So $v_{\mathfrak{p}^{\prime}}(xy-1)\geq 1$. Hence 1xy1-xy is not invertible in $R$, which contradicts to $x\in J(R)$.
In conclusion we have $J(R)=\bigcap\limits_{\mathfrak{p}\in\Delta}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}.$

### v3

@@ 2.2. Global class field theory
In this section, we recall some standard results from global class field theory. We begin by introducing the notion of a modulus, which generalizes the role of an integer $m$ in classical modular arithmetic.
@@
$I_{\mathfrak{m}}\coloneqq\langle\mathfrak{p}\subseteq\mathcal{O}_{K}:\mathfrak{m}(\mathfrak{p})=0\rangle$
@@
$(\alpha)\coloneqq\alpha\mathcal{O}_{K}\;\in\;I_{K}.$
@@
$P_{\mathfrak{m}}\coloneqq i(K_{\mathfrak{m},1})\;\leq\;I_{\mathfrak{m}}.$
@@
$C_{\mathfrak{m}}\coloneqq I_{\mathfrak{m}}/P_{\mathfrak{m}}.$
@@
$\psi_{L/K}(\mathfrak{a})\;\coloneqq\;\prod(\mathfrak{p}_{i},L/K)^{e_{i}}.$
@@
where the first product runs over all finite primes $\mathfrak{P}$ of $L$ above $\mathfrak{p}$, and the second product runs over all real embeddings $\Omega$ of $L$ extending the real embedding $\omega$ of $K$.
@@ 3.1. Notations and Some Examples
$S(A/K)\coloneqq\{\textup{Trd}(x):x\in A,\ \textup{Nrd}(x)=1\},\qquad T(A/K)\coloneqq S(A/K)+S(A/K),$
@@
$H_{a,b}\coloneqq K\oplus\alpha K\oplus\beta K\oplus\alpha\beta K,$
where $\alpha^{2}=a$, $\beta^{2}=b$, $\alpha\beta=\beta\alpha$.
@@
$\mathcal{O}_{\mathfrak{p}}\coloneqq(\mathcal{O}_{K})_{\mathfrak{p}}=R_{\mathfrak{p}}\cap K$
@@
If $A/K$ splits at all real places of $K$ (this holds if $K$ is a global function field, or if $K$ is a number field and $a,b\in K^{\times}$ are totally positive), then
@@
$T_{a,b}$ is a diophantine set over $K$ by definition. This fact forms the basis for proving that many other sets are diophantine in [EM18, DIT18, PAR13]. Moreover, $T_{a,b}$ is uniformly existentially defined in $a$ and $b$, since we have the following formula for $S_{a,b}$:
$S_{a,b}=\{2x_{1}\in K:\exists x_{2},x_{3},x_{4}\text{ with }x_{1}^{2}-ax_{2}^{2}-bx_{3}^{2}+abx_{4}^{2}=1\},$
together with the definition $T_{a,b}=S_{a,b}+S_{a,b}$.
The following definition appears in Definition 9 of [KOE16] for $K=\mathbb{Q}$, and is also given after the proof of Lemma 3.8 in [PAR13] as well as after the proof of Lemma 3.3 in [EM18].
@@
Let $L\coloneqq K(\sqrt{a},\sqrt{b})/K$ be an abelian extension with Galois group isomorphic to $(\mathbb{Z}/2\mathbb{Z})^{2}$, and let $\psi_{L/K}:I_{\mathfrak{m}}\rightarrow\textup{Gal}(L/K)$ be the global Artin map with admissible modulus $\mathfrak{m}$. We define the following notations:
  • $\mathbb{P}(p)\coloneqq\{\mathfrak{p}\in S_{K}^{f}:v_{\mathfrak{p}}(p)\text{ is odd}\}$.
  • For $(i,j)\in\textup{Gal}(L/K)$, $i,j\in\{\pm 1\}$, we define
    $\mathbb{P}^{(i,j)}\coloneqq\{\mathfrak{p}\in S_{K}^{f}:\mathfrak{p}\in I_{\mathfrak{m}}\text{ and }\psi_{L/K}(\mathfrak{p})=(i,j)\}.$
  • $\mathbb{P}^{(i,j)}(p)\coloneqq\mathbb{P}(p)\cap\mathbb{P}^{(i,j)}$.
Here, $(1,1)$ is the identity in the Galois group. (1,1)(-1,1) is the element which sends $\sqrt{a}$ to $\sqrt{a}$ and fixes $\sqrt{b}$. (1,1)(1,-1) is the element which sends $\sqrt{b}$ to $\sqrt{b}$ and fixes $\sqrt{a}$. Finally, (1,1)(-1,-1) is the element which sends $\sqrt{a}$ to $\sqrt{a}$ and $\sqrt{b}$ to $\sqrt{b}$.
@@
(1) If $K$ is a number field, there exist $a,b\in K^{\times}$ with disjoint supports and $\textup{Gal}(K(\sqrt{a},\sqrt{b})/K)\cong(\mathbb{Z}/2\mathbb{Z})^{2}$ such that for any $p\in K^{\times}$ coprime to the admissible modulus $\mathfrak{m}$, we have:
@@
(2) If $K$ is a global function field with $\operatorname{char}(K)\neq 2$, there exist $a,b,c,d\in K^{\times}$ such that $a$ and $b$ have disjoint supports and $\textup{Gal}(K(\sqrt{a},\sqrt{b})/K)\cong(\mathbb{Z}/2\mathbb{Z})^{2}$. For any $p\in K^{\times}$ coprime to the admissible modulus $\mathfrak{m}$, we have:
@@
We first address (1), which was proved in [PAR13]. By Lemma 3.9 of [PAR13], for any $p$ coprime to $\mathfrak{m}$, the sets on the left and right hand sides of each equation differ at most by the primes dividing the modulus. Then, Lemma 3.19 and Lemma 3.20 in [PAR13] provide a careful choice of $a$ and $b$ such that these sets are equal.
Part (2) follows from Corollary 3.11 in [EM18].
The fact that $a$ and $b$ have disjoint supports is guaranteed by Lemma 3.19 (4) of [PAR13] and Lemma 3.8 (4) of [EM18]. ∎
From now on, unless otherwise stated, the parameters $a,b$ (or $a,b,c,d$) are as fixed in Lemma 3.4, and $L$ denotes the field $K(\sqrt{a},\sqrt{b})$, depending on whether $K$ is a number field or a function field.
@@
Let $K$ be a global field with $\operatorname{char}(K)\neq 2$, and keep the constants $a,b,c,d$ as above. For each $\sigma\in\textup{Gal}(L/K)$, we define the ring $R_{p}^{\sigma}$ as follows:
  1. (a)
    If $K$ is a number field, then:
    • $R_{p}^{(-1,-1)}\coloneqq\bigcap_{\mathfrak{p}\in\Delta_{a,p}\cap\Delta_{b,p}}\mathcal{O}_{\mathfrak{p}}$
    • $R_{p}^{(1,-1)}\coloneqq\bigcap_{\mathfrak{p}\in\Delta_{ab,p}\cap\Delta_{b,p}}\mathcal{O}_{\mathfrak{p}}$
    • $R_{p}^{(-1,1)}\coloneqq\bigcap_{\mathfrak{p}\in\Delta_{a,p}\cap\Delta_{ab,p}}\mathcal{O}_{\mathfrak{p}}$
    • $R_{p,q}^{(1,1)}\coloneqq\bigcap_{\mathfrak{p}\in\Delta_{ap,q}\cap\Delta_{bp,q}}\mathcal{O}_{\mathfrak{p}}$
  2. (b)
    If $K$ is a global function field, then:
    • $R_{p}^{(-1,-1)}\coloneqq\bigcap_{\mathfrak{p}\in\Delta_{a,p}\cap\Delta_{b,p}}\mathcal{O}_{\mathfrak{p}}$
    • $R_{p}^{(1,-1)}\coloneqq\bigcap_{\mathfrak{p}\in\Delta_{ab,p}\cap\Delta_{b,p}\cap\Delta_{a,cp}}\mathcal{O}_{\mathfrak{p}}$
    • $R_{p}^{(-1,1)}\coloneqq\bigcap_{\mathfrak{p}\in\Delta_{a,p}\cap\Delta_{ab,p}\cap\Delta_{b,dp}}\mathcal{O}_{\mathfrak{p}}$
    • $R_{p,q}^{(1,1)}\coloneqq\bigcap_{\mathfrak{p}\in\Delta_{ap,q}\cap\Delta_{bp,q}}\mathcal{O}_{\mathfrak{p}}$
For $K=\mathbb{Q}$, the definitions of $R_{p}^{\sigma}$ and $R_{p,q}^{(1,1)}$ above coincide with Definition 7 in [KOE16]. Their generalizations to number fields and global fields are given in Definition 3.10 of [PAR13] and Definition 3.12 of [EM18], respectively.
The following lemma is well known (see Lemma 3.2.6 and Theorem 3.2.7 in [EP05]).
Lemma 3.6.
Let $\Delta$ be a finite subset of $S_{K}^{f}$. Recall that $J(R)$ denotes the Jacobson radical of a ring $R$. If
$R=\bigcap_{\mathfrak{p}\in\Delta}\mathcal{O}_{\mathfrak{p}},$
then
$J(R)=\bigcap_{\mathfrak{p}\in\Delta}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}.$
The following lemma describes the behavior of addition for finite intersections of valuation rings and their Jacobson radicals.
Lemma 3.7.
Let $\Delta_{1},\Delta_{2}$ be two finite subsets of $S_{K}^{f}$. Then:
  • (1)
    $\bigcap_{\mathfrak{p}\in\Delta_{1}}\mathcal{O}_{\mathfrak{p}}+\bigcap_{\mathfrak{p}\in\Delta_{2}}\mathcal{O}_{\mathfrak{p}}=\bigcap_{\mathfrak{p}\in\Delta_{1}\cap\Delta_{2}}\mathcal{O}_{\mathfrak{p}}.$
  • (2)
    $\bigcap_{\mathfrak{p}\in\Delta_{1}}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}+\bigcap_{\mathfrak{p}\in\Delta_{2}}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}=\bigcap_{\mathfrak{p}\in\Delta_{1}\cap\Delta_{2}}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}.$
Proof.
For both (1) and (2), the left-hand sides are obviously contained in the right-hand sides. For the reverse inclusion, let $t\in\{0,1\}$ and $x\in K$ be such that $v_{\mathfrak{p}}(x)\geq t$ for all $\mathfrak{p}\in\Delta_{1}\cap\Delta_{2}$.
By weak approximation, there exists $x_{1}\in K$ such that
$v_{\mathfrak{p}}(x_{1})\geq t\quad\text{for all }\mathfrak{p}\in\Delta_{1},$
and
$v_{\mathfrak{p}}(x_{1}-x)\geq t\quad\text{for all }\mathfrak{p}\in\Delta_{2}\setminus\Delta_{1}.$
Set $x_{2}=x-x_{1}$. Then, for all $\mathfrak{p}\in\Delta_{1}\cap\Delta_{2}$, we have
$v_{\mathfrak{p}}(x_{2})=v_{\mathfrak{p}}(x-x_{1})\geq\min\{v_{\mathfrak{p}}(x),v_{\mathfrak{p}}(x_{1})\}\geq t.$
Thus $x=x_{1}+x_{2}$, which is of the form appearing on the left-hand side. ∎
The following definition summarizes Definition 3.21 and Definition 3.24 in [PAR13], which also appear as Definition 3.13 and Definition 3.16 in [EM18].
Definition 3.8.
For each $\sigma\in\textup{Gal}(L/K)$, define
$\Phi_{\sigma}\coloneqq\{p\in K^{\times}:(p)\in I_{\mathfrak{m}},\ \psi_{L/K}((p))=\sigma,\textup{ and }\mathbb{P}(p)\subset\mathbb{P}^{(1,1)}\cup\mathbb{P}^{\sigma}\},$
$\widetilde{\Phi_{\sigma}}\coloneqq{K^{\times}}^{2}\cdot\Phi_{\sigma},$
and
$\Psi_{K}\coloneqq\Big\{(p,q)\in\widetilde{\Phi_{(1,1)}}\times\widetilde{\Phi_{(-1,-1)}}\ \Big|\ \prod_{\mathfrak{p}\mid\mathfrak{m}}(ap,q)_{\mathfrak{p}}=-1\textup{ and }p\in a\cdot{K^{\times}}^{2}\cdot(1+J(R_{q}^{(-1,-1)}))\Big\}.$
For later use, we summarize several results from [EM18] and [PAR13] as follows.
@@
Let $K$ be a global field with $\operatorname{char}(K)\neq 2$, and keep the notations introduced above. Then the following statements hold:
  • (1)
    The sets $\Phi_{\sigma}$, $\Psi_{K}$, $R_{p}^{\sigma}$, $J(R_{p}^{\sigma})$, $R_{p,q}^{(1,1)}$, and $J(R_{p,q}^{(1,1)})$ are all diophantine sets over $K$.
  • (2)
    For any $p\in\Phi_{\sigma}$ and any $\sigma\in\textup{Gal}(L/K)$ with $\sigma\neq(1,1)$, the set $\mathbb{P}^{\sigma}(p)$ is nonempty.
  • (3)
    Let $\sigma\in\textup{Gal}(L/K)$ with $\sigma\neq(1,1)$, and let $\mathfrak{p}\nmid\mathfrak{m}$ be a prime of $K$ satisfying $\psi_{L/K}(\mathfrak{p})=\sigma$. Then there exists an element $p\in\Phi_{\sigma}$ such that
    $\mathbb{P}^{\sigma}(p)=\{\mathfrak{p}\}.$
  • (4)
    For any $(p,q)\in\Psi_{K}$, we have
    $\emptyset\neq\Delta_{ap,q}\cap\Delta_{bp,q}\subset I_{\mathfrak{m}}.$
  • (5)
    For each prime ideal $\mathfrak{p}_{0}$ satisfying $\mathfrak{p}_{0}\nmid\mathfrak{m}$ and $\psi_{L/K}(\mathfrak{p}_{0})=(1,1)$, there exists $(p,q)\in\Psi_{K}$ such that
    $\Delta_{ap,q}\cap\Delta_{bp,q}=\{\mathfrak{p}_{0}\}.$
    Moreover, we may require that $v_{\mathfrak{p}_{0}}(p)=1$, $v_{\mathfrak{p}_{0}}(q)=0$, and that $q$ is not a square modulo $\mathfrak{p}_{0}$.
@@
This proposition follows from Lemma 3.14 and Lemma 3.17 in [EM18] for global function fields, and from Lemma 3.22 and Lemma 3.25 in [PAR13] for number fields. Only the last assertion in (5) requires additional justification, which we provide below.
Number field case. The proof of Lemma 3.25(c) in [PAR13] is based on an explicit construction of the pair $(p,q)$. In the choice of $p,q\in K$, condition (B) for $q$ requires that the Legendre symbol $\big(\frac{q}{\mathfrak{p}_{0}}\big)$ is equal to 1-1. This implies that $v_{\mathfrak{p}_{0}}(q)=0$ and that $q$ is not a square modulo $\mathfrak{p}_{0}$.
Condition (4) for $p$ is $(q,p)_{\mathfrak{p}_{0}}=-1$. Using the explicit formula for the Hilbert symbol, we compute
$-1=(q,p)_{\mathfrak{p}_{0}}$ $=(-1)^{v_{\mathfrak{p}_{0}}(q)v_{\mathfrak{p}_{0}}(p)\frac{|\mathbb{F}_{\mathfrak{p}_{0}}|-1}{2}}\big(\tfrac{q}{\mathfrak{p}_{0}}\big)^{v_{\mathfrak{p}_{0}}(p)}\big(\tfrac{p}{\mathfrak{p}_{0}}\big)^{v_{\mathfrak{p}_{0}}(q)}$
$=\big(\tfrac{q}{\mathfrak{p}_{0}}\big)^{v_{\mathfrak{p}_{0}}(p)},$
since $v_{\mathfrak{p}_{0}}(q)=0$. As $\big(\tfrac{q}{\mathfrak{p}_{0}}\big)=-1$, this implies that $v_{\mathfrak{p}_{0}}(p)$ is odd. By multiplying $p$ by a square in $K^{\times}$, we may further assume that $v_{\mathfrak{p}_{0}}(p)=1$.
Global function field case. This is addressed in the proof of Lemma 3.17(3) in [EM18]. In the construction, the element $q$ again satisfies $\big(\tfrac{q}{\mathfrak{p}_{0}}\big)=-1$. Since $\mathfrak{p}_{0}\in\Delta_{ap,q}$, we have $(ap,q)_{\mathfrak{p}_{0}}=-1$. By computing the Hilbert symbol and using the fact that $\mathfrak{p}_{0}\nmid a$, we conclude that $v_{\mathfrak{p}_{0}}(p)$ is odd. This is stated explicitly in the third-to-last line of the proof of Theorem 1.3 in [EM18]. ∎
Remark 3.10.
In fact, the sets $R_{p}^{\sigma}$, $J(R_{p}^{\sigma})$, $R_{p,q}^{(1,1)}$, and $J(R_{p,q}^{(1,1)})$ are not only diophantine, but are defined uniformly in the parameters $p$ and $q$. More precisely, the following sets are diophantine:
$\{(p,x)\in K^{\times}\times K:x\in R_{p}^{\sigma}\},$
$\{(p,x)\in K^{\times}\times K:x\in J(R_{p}^{\sigma})\},$
$\{(p,q,x)\in K^{\times}\times K^{\times}\times K:x\in R_{p,q}^{(1,1)}\},$
$\{(p,q,x)\in K^{\times}\times K^{\times}\times K:x\in J(R_{p,q}^{(1,1)})\}.$
For the rings $R_{p}^{\sigma}$ and $R_{p,q}^{(1,1)}$, this uniformity follows from the uniform definability of $T_{a,b}$ (see Remark 3.2), together with Lemma 3.1 and Lemma 3.7(1). For the corresponding Jacobson radicals, the uniformity follows from Lemma 3.17 of [PAR13], Lemma 3.6, and Lemma 3.7(2).
3.2. Diagonalization of quadratic forms
Since we always work over a global field $K$ with characteristic not equal to $2$, we obtain the following proposition concerning the diagonalization of quadratic forms.